OpenStudy (anonymous):

if i have a half dollar and is 8X face value how much will it be worth

OpenStudy (anonymous):

\[8\times .5=4\] so $4

OpenStudy (anonymous):

ok thanks

OpenStudy (ash2326):

\[ 8\times \frac{1}{2}=$4\]